OpenStudy (anonymous):

Show that \[\int_0^\infty \frac{\ln x}{1+x^2}\,dx=0\]

OpenStudy (anonymous):

*

OpenStudy (anonymous):

I have a partial solution, but I'm interested in seeing if there are other exact methods...

OpenStudy (anonymous):

What I did was introduce a parameter \(b\) with \[I(b)=\int_0^\infty \frac{\ln bx}{1+x^2}\,dx\] and differentiating yields \[I'(b)=\int_0^\infty \frac{dx}{1+x^2}=\frac{\pi}{2b}\] Integrating gives \[I(b)=\frac{\pi}{2}\ln|b|+C\] and so setting \(b=1\) gives the desired result, but I don't see how to establish that \(C\) must also vanish.

OpenStudy (zarkon):

just use a substitution (\(u=\ln(x)\))

OpenStudy (zarkon):

\[\int\limits_0^\infty \frac{\ln x}{1+x^2}~dx=\int\limits_{-\infty}^{\infty}\frac{u}{e^{-u}+e^{u}}~du\]

OpenStudy (zarkon):

\[\frac{u}{e^{-u}+e^{u}}\] is odd

OpenStudy (anonymous):

Then integrate by parts... or yes, use symmetry.

ganeshie8 (ganeshie8):

nice :)

OpenStudy (anonymous):

As an aside, as far as symmetry goes, how does odd-ness here help us conclude the integral is zero, whereas for something like \(\int x^3~dx\) we have a divergent integral?

OpenStudy (anonymous):

Oh never mind, I got it...

OpenStudy (zarkon):

\[\int\limits_{0}^{\infty}\frac{u}{e^-{u}+e^{u}}du\] clearly converges thus \[\int\limits_{-\infty}^{\infty}\frac{u}{e^{-{u}}+e^{u}}du=\lim_{a\to\infty}\int\limits_{-a}^{a}\frac{u}{e^{-{u}}+e^{u}}du\]

OpenStudy (anonymous):

Right, I had the wrong image in mind... Thanks for the (simpler) approach. If anyone has a suggestion or correction for the DUIS approach, by all means do let me know. I was hoping to practice something new and this integral seemed like a good candidate.

ganeshie8 (ganeshie8):

a modest attempt would be to try \(x = \tan u\)

ganeshie8 (ganeshie8):

\[\begin{align}\int\limits_0^\infty \frac{\ln x}{1+x^2}~dx ~&\stackrel{x=\tan u}{=}~\int\limits_0^{\pi/2} \ln (\tan u) du\\~\\&~~ = -\int\limits_0^{\pi/2} \ln (\tan u) du\\~\\ &\implies 0\end{align}\]

OpenStudy (anonymous):

Nice, I also had arrived at a similar result with IBP showing that \(\int=-\int\).

OpenStudy (jhannybean):

*

OpenStudy (anonymous):

If I'm remembering correctly, @Kainui might have some expertise on the DUIS matter...